Help...
State whether the conclusion Is correct and why or why not
-2(a-3)=6
-2a-6=6

Answer by stanbon(75887)

State whether the conclusion Is correct and why or why not
-2(a-3)=6
-2a-6=6
------
Ans:: not correct.
Improper application of the distributive law::
-2(a-3) = 6
-2a + 6 = 6 is the proper application of the distributive law.
-------
Cheers,
Stan H.
-------